# Intelligent Surfaces Empowered Wireless Network: Recent Advances and The Road to 6G

### 15 Jan 2024 v2

@@ I-A Overview of 6G
The global mobile traffic in the next decade is envisaged to increase explosively and reach over 5000 exabytes per month in 2030 [1]. Such a demand will pose a huge challenge to the current fifth-generation (5G) wireless network in terms of energy consumption, cost, complexity, etc. To address this challenge and foster new applications, worldwide academic and industrial communities, together with standardization organizations, have been devoting their efforts to embrace the next/sixth-generation (6G) wireless network in the future. Starting from the first global 6G research program named “6Genesis” in 2018 [2], extensive preliminary projects, works, and standard proposals have been pursued to define/identity the 6G framework, usage scenarios, key technical requirements, and enabling technologies. A very recent significant progress is that in June 2023, a new and complete 6G vision has been proposed by the International Mobile Telecommunications for 2030 and beyond (IMT-2030) promotion group during the 44th meeting of the international telecommunication union radiocommunication sector working party 5D (ITU-R WP5D) [3, 4]. Evolving from 5G, 6G is envisioned to provide global seamless coverage, expand the utilized spectrum resources beyond the millimeter wave (mmWave) frequency bands, and establish the physical-digital world mapping to support a wide range of new applications and services intelligently, which will drive the next wave of digital economic growth, as well as sustainable far-reaching societal changes and digital equality.
@@ I-B1 Historical Development of ISs
Research on IS can be traced back to the theoretical works studying the interaction between electromagnetic (EM) waves and metamaterials. As shown in Fig. 2, the transition conditions for the average EM fields across a surface with a type of metamaterial called metafilm were studied in 2003, showing that the EM wave can be controlled by electrically changing the properties of a metafilm [6]. Subsequently, an intelligent wall equipped with a certain type of metamaterial called active frequency selective surface was proposed to manipulate the propagation environment inside a building [7]. Afterwards, the controllability of metamaterials over EM waves underwent further development. In 2014, tunable metasurfaces were developed to construct a spatial microwave binary phase modulator [8]. Meanwhile, programmable metamaterials were proposed to manipulate EM waves by digitally coding the metamaterials with a field-programmable gate array (FPGA) [9], which greatly extended the flexibility of controlling the EM wave.
@@
The aforementioned development of programmable metamaterials/surfaces promoted the formation of the new IS concept and its applications in wireless communications. In 2018, the large intelligent surface (LIS) was proposed for enhancing wireless sensing and communication performance in [10, 11, 12]. As an extension of the conventional massive multiple-input-multiple-output (MIMO) technology, LIS, like massive MIMO, is also based on the active array architecture, which incurs higher hardware cost and power consumption with increasingly more antenna elements employed. In the same year, a software-controlled programmable metamaterial architecture named HyperSurface Tile was proposed in [13] for proactively controlling the radio propagation environment to enhance the wireless communication performance. In addition, a similar concept called intelligent reflecting surface (IRS) was proposed independently in [14], which, for the first time, investigated the joint active and passive beamforming design problem in an IRS-aided wireless system for maximizing the communication spectral efficiency (SE). In particular, a quadratic power scaling law with the increasing number of IRS elements was revealed in [14], which unveiled the more cost-effective performance gain of IRS passive beamforming as compared to the conventional active beamforming (via e.g., massive MIMO and LIS) with only a linear power gain over the number of antenna elements. Whereafter, other similar IS structures, e.g., reconfigurable intelligent surface (RIS) [15] and holographic MIMO surface [16] were also proposed to extend the metric/structure of IRS from SE boost/discrete-element implementation to energy efficiency (EE) improvement and continuous aperture, respectively. To extend the signal coverage of reflection-based IRS, the intelligent omni-surface (IOS) [17] and the Simultaneous Transmitting and Reflecting RIS (STAR-RIS) [18] were subsequently proposed to enable both signal reflection and refraction of ISs. A more recent IS development was the general Beyond Diagonal RIS (BD-RIS) architecture [19], which includes the above-mentioned IS architectures as special cases by enabling more sophisticated signal processing across IS elements.
@@ I-B2 Classifications of ISs
  • Active/Passive ISs: The passive IS only reflects/refracts the incident signal with controllable phase shift and/or amplitude without amplification. In contrast, the active IS is able to amplify the impinging signal while manipulating its phase shift at the same time.
@@ I-B3 Applications of ISs in 6G
As shown in Fig. 3, ISs are envisioned to facilitate and enable a wide range of applications for future 6G usage scenarios. For example, to enable immersive communication applications, such as extended reality (XR) and ultra-high definition (UHD) video services usually supported by high-frequency wideband signal transmission, ISs can help overcome the severe path loss and bypass obstacles by providing high beamforming gains and additional propagation paths, respectively. In addition, to facilitate hyper-reliable and low-latency communication applications, such as industrial control, telemedicine and emergency services, ISs can be utilized to improve the channel quality and transmission reliability by creating the dominant line-of-sight (LoS) path to mitigate the undesired multipath fading. Particularly, for services with stringent requirements on reliability and latency, especially in high-mobility scenarios such as fully autonomous driving, ISs can be either deployed at the roadside to enhance the signal strength for the vehicles (or passengers inside vehicles) on the road, or mounted on the vehicles to mitigate the fast channel fading arising from their high mobility [20]. Moreover, to support massive communications in e.g., smart cities, environmental monitoring and other IoT applications, ISs can be used to improve the system throughput and fairness for densely deployed devices by jointly designing reflections and their non-orthogonal multiple access (NOMA) scheme [21]. Moreover, refraction-based ISs can be deployed to increase the number of connections with devices by extending their signal coverage.
@@
Furthermore, apart from the aforementioned three scenarios, ISs can also effectively facilitate applications in the three newly introduced scenarios for 6G. For instance, to promote AI and communication applications such as distributed computing and digital twins, large-aperture ISs can be employed to meet the requirement on high data rates and quality of service (QoS) by providing large passive beamforming gains; while AI methods like deep learning can also be utilized to perform signal detection, modulation, channel estimation, beamforming/reflection, and resource allocation at the IS side, which are expected to achieve better performance than traditional signal processing and optimization methods [22]. Besides, to support the ubiquitous connectivity vision, ISs can be deployed on mobile platforms such as unmanned aerial vehicles (UAVs) or balloons to extend the coverage, especially for scarcely covered areas. Apart from the above communication applications, ISs can also be employed to improve both the sensing and communication performance in ISAC applications. For instance, to enable ISAC services such as navigation, high-precision positioning, user activity detection and tracking, ISs can be flexibly deployed on the surface of walls, buildings, and vehicles as additional reference nodes to enhance the positioning accuracy of the future dual-functional radar and communication (DFRC) base station (BS) [23, 24], as well as mounted on targets like automobiles or UAVs to enlarge their radar cross section (RCS) for high-resolution detection and highly accurate real-time tracking [20, 24]. In addition to the above, ISs can assist in many other applications, such as enhancing the physical-layer security [25] and boosting both the communication rate and the energy transmission efficiency for simultaneous wireless information and power transfer (SWIPT) [26].
@@ I-C Standardization Activities
As an innovative radio access technology, IS has already been included in the ITU-R IMT-2030 report on future technology trends, as a candidate technology for enhancing the radio interface of future telecommunication standards [27]. During the last two years, IS has been under evaluation for its inclusion in future telecommunication standards, and pre-standarization and standardization bodies are currently working on it. In this subsection, we briefly overview the most relevant standarization activities on IS. Also, we put them in context with other standarization activities focusing on technologies that have strong ties with IS, notably terahertz (THz) communications and ISAC.
@@ I-C1 Third Generation Partnership Program (3GPP)
  • NR smart repeaters from China Mobile Communications Corporation (CMCC) (RWS-210339) – Study candidate side control information such as timing, beam, and bandwidth information for smart repeaters (RAN1, RAN4); and specification of radio frequency requirements considering side control information (RAN4).
@@
  • NR smart repeaters for Rel-18 from Qualcomm (RWS-210019) – Work in Rel-18 on smart repeaters, including the study of side control information (RAN1), protocol support (RAN2), and radio frequency and electromagnetic compatibility requirements (RAN4).
@@
As far as 3GPP is concerned, Rel-18 will include, on the other hand, a study item (RP-213700 [29], TR-38.867 [30]) on network controlled repeaters (NCRs). These repeaters are intended to address the limitations of conventional (radio frequency) repeaters as well as to provide enhanced low-cost coverage. Conventional amplify-and-forward repeaters are considered which operate with minimal processing and do not require signal decoding and reconstruction. Therefore, they operate at a lower complexity and latency than decode-and-forward repeaters. The price to pay for the reduction in complexity lies in a reduced link quality and increased network interference, which is due to the interference/noise amplification. When applied to high frequency bands (above 24 GHz), amplify-and-forward repeaters provide limited coverage. In fact, static directional repeaters cannot adapt to the channel conditions and user mobility, which is necessary in communication systems where directional communications and high beamforming gains are necessary. NCRs will be designed to include an in-band control interface to configure and manage their behavior, with the objective of reducing the amplification of noise and supporting communications with a better spatial directivity. A detailed description of NCRs in the context of 3GPP standardization can be found in [31]. NCRs are often considered a stepping stone for ISs, since the control channel may be reused. Recent research works have also compared NCRs and ISs, as summarized in [31].
@@ I-C2 European Telecommunication Standards Institute (ETSI)
TABLE I: List of Representative Overview/Survey/Tutorial Papers on ISs
@@
Debunk three myths of the RIS and raised two critical questions to inspire the future research.
@@
2021
@@ I-D Aims and Organization
In contrast to the existing overview, survey, and tutorial papers on ISs listed in Table I-C2, this paper aims to provide a comprehensive overview of the latest development and advancement of ISs, with an emphasis on recent research results as well as new IS architectures and their design issues. Specifically, this paper first provides an in-depth overview of the state-of-the-art results on the design of the commonly adopted reflection-based IS including its reflection optimization, deployment, signal modulation, wireless sensing, and ISAC. Next, we delve into the recent progress in advanced IS architectures and their associated new design issues and solutions. Moreover, we highlight the important problems that remain unsolved to motivate future research on ISs. We hope that this paper will help unveil the immense potential of ISs for applications in future wireless networks such as 6G, and provide useful and up-to-date guidance for the future research and development for ISs.

### 24 Mar 2024 v3

@@ I-A Overview of 6G
The global mobile traffic in the next decade is envisaged to increase explosively and reach over 5000 exabytes per month in 2030 [1]. Such a demand will pose a huge challenge to the current fifth-generation (5G) wireless network in terms of energy consumption, cost, complexity, etc. To address this challenge and foster new applications, worldwide academic and industrial communities, together with standardization organizations, have been devoting their efforts to embrace the next/sixth-generation (6G) wireless network in the future. Starting from the first global 6G research program named “6Genesis” in 2018 [2], extensive preliminary projects, works, and standard proposals have been pursued to define/identity the 6G framework, usage scenarios, key technical requirements, and enabling technologies [3]. A very recent significant progress is that in June 2023, a new and complete 6G vision has been proposed by the International Mobile Telecommunications for 2030 and beyond (IMT-2030) promotion group during the 44th meeting of the international telecommunication union radiocommunication sector working party 5D (ITU-R WP5D) [4, 5]. Evolving from 5G, 6G is envisioned to provide global seamless coverage, expand the utilized spectrum resources beyond the millimeter wave (mmWave) frequency bands, and establish the physical-digital world mapping to support a wide range of new applications and services intelligently, which will drive the next wave of digital economic growth, as well as sustainable far-reaching societal changes and digital equality.
@@ I-B1 Historical Development of ISs
As shown in Fig. 2, ISs can be traced back to the early study on a special type of antenna array called reflectarray [7]. Reflectarray operates with an active source radiating radio-frequency (RF) signals towards a nearby reflective surface formed by arrays of passive antennas to achieve desired radiation patterns, which was aimed at reducing the hardware cost of conventional active antenna arrays. Subsequently, the research on investigating the more sophisticated interaction between electromagnetic (EM) waves and metamaterials was also pursued (e.g., [8, 9]). In particular, the transition conditions for the average EM fields across a surface with a type of metamaterial called metafilm were studied in 2003, showing that the EM wave can be controlled by electrically changing the properties of a metafilm [10]. This result inspired the new idea of deploying metamaterial-made IS in the environment to alter wireless signal propagation, which thus opened up a new direction for IS beyond the traditional reflectarray. Subsequently, an intelligent wall equipped with a certain type of metamaterial called active frequency selective surface was proposed to manipulate the propagation environment inside a building [11]. Afterwards, the controllability of metamaterials over EM waves underwent further development. In 2014, tunable metasurfaces were developed to construct a spatial microwave binary phase modulator [12]. Meanwhile, programmable metamaterials were proposed to manipulate EM waves by digitally coding the metamaterials with a field-programmable gate array (FPGA) [13], which greatly extended the flexibility of controlling the EM wave.
@@
The aforementioned development of programmable metamaterials/surfaces promoted the formation of the new IS concept and its applications in wireless communications. In 2018, the large intelligent surface (LIS) was proposed for enhancing wireless sensing and communication performance in [14, 15, 16]. As an extension of the conventional massive multiple-input-multiple-output (MIMO) technology, LIS, like massive MIMO, is also based on the active array architecture, which incurs higher hardware cost and power consumption with increasingly more antenna elements employed. In the same year, a software-controlled programmable metamaterial architecture named HyperSurface Tile was proposed in [17] for proactively controlling the radio propagation environment to enhance the wireless communication performance. In addition, a similar concept called intelligent reflecting surface (IRS) was proposed independently in [18], which, for the first time, investigated the joint active and passive beamforming design problem in an IRS-aided wireless system for maximizing the communication spectral efficiency (SE). In particular, a quadratic power scaling law with the increasing number of IRS elements was revealed in [18], which unveiled the more cost-effective performance gain of IRS passive beamforming as compared to the conventional active beamforming (via e.g., massive MIMO and LIS) with only a linear power gain over the number of antenna elements. IS structures that are not limited to reflection, i.e., reconfigurable intelligent surface (RIS) were discussed in [19] and [20]. Also, holographic MIMO surface [21] was proposed to encompass relay-type and transmitter-type, as well as discrete and continuous implementations. To extend the signal coverage of reflection-based IRS, the intelligent omni-surface (IOS) [22] and the Simultaneous Transmitting and Reflecting RIS (STAR-RIS) [23] were subsequently proposed to enable both signal reflection and refraction of ISs. A more recent IS development was the general Beyond Diagonal RIS (BD-RIS) architecture [24], which includes the above-mentioned IS architectures as special cases by enabling more sophisticated signal processing across IS elements. Another architecture based on multi-layer ISs is refereed to as stacked intelligent metasurface [25]. ISs have put forth the emerging concept of smart radio environment that can provide uninterrupted wireless connectivity, in an energy sustainable manner by recycling existing radio waves [26, 27, 28].
@@ I-B2 Classifications of ISs
  • Active/Passive ISs: The passive IS only reflects/refracts the incident signal with controllable phase shift and/or amplitude without amplification. In contrast, the active IS is able to amplify the impinging signal while manipulating its phase shift at the same time. [^2] In the rest of this paper, the term passive refers to ISs with no power amplification capabilities.
@@ I-B3 Applications of ISs in 6G
As shown in Fig. 3, ISs are envisioned to facilitate and enable a wide range of applications for future 6G usage scenarios. For example, to enable immersive communication applications, such as extended reality (XR) and ultra-high definition (UHD) video services usually supported by high-frequency wideband signal transmission, ISs can help overcome the severe path loss and bypass obstacles by providing high beamforming gains and additional propagation paths, respectively [29]. Additionally, to facilitate hyper-reliable and low-latency communication applications, such as industrial control, telemedicine and emergency services, ISs can be utilized to improve the channel quality and transmission reliability by creating the dominant line-of-sight (LoS) path to mitigate the undesired multipath fading. Particularly, for services with stringent requirements on reliability and latency, especially in high-mobility scenarios such as fully autonomous driving, ISs can be either deployed at the roadside to enhance the signal strength for the vehicles (or passengers inside vehicles) on the road, or mounted on the vehicles to mitigate the fast channel fading arising from their high mobility [30]. Moreover, to support massive communications in e.g., smart cities, environmental monitoring and other IoT applications, ISs can be used to improve the system throughput and fairness for densely deployed devices by jointly designing reflections/refractions and their non-orthogonal multiple access (NOMA) scheme [31, 32].
In addition to the aforementioned three scenarios, ISs can effectively facilitate applications in the newly introduced scenarios for 6G. Under the integrated AI and communication scenario, especially the edge AI scenario where edge servers aggregate data from mobile devices over wireless links to perform AI tasks such as federated learning, ISs can be utilized to provide favorable wireless channel environment for real-time data aggregation and exchange, distributed computing, remote monitoring and controlling, etc., thereby improving AI performance and scalability [33, 34]. On the other hand, AI algorithms can also be promisingly applied to solve non-convex, highly nonlinear and large-scale design and optimization problems in IS-aided wireless networks, such as passive beamforming, channel estimation, optimal deployment and resource allocation [35]. Besides, the use of massive and flexibly deployed ISs in proper locations is appealing to achieve the ubiquitous connectivity even in a complex propagation environment with rich scatterers/obstacles. In addition, ISs can be deployed on maritime or aerial mobile platforms (e.g., ships, unmanned aerial vehicles (UAVs), balloons, airships, satellites, and so on) to provide wireless coverage for scarcely covered areas [36, 37, 38]. Furthermore, the flexibility of ISs to reflect or refract wireless signals enables coverage for dead zones in conventional terrestrial networks [39].
@@
Apart from communication applications, ISs can also be employed to improve both the sensing and communication performance in ISAC applications. For instance, to enable ISAC services such as navigation, high-precision positioning, user activity detection and tracking, ISs can be flexibly deployed on the surface of walls, buildings, and vehicles as additional reference nodes to enhance the positioning accuracy of the future dual-functional radar and communication (DFRC) BS [40, 41], as well as mounted on targets like automobiles or UAVs to enlarge their radar cross section (RCS) for high-resolution detection and highly accurate real-time tracking [30, 41]. In addition to the above, ISs can assist in many other applications, such as enhancing the physical-layer security [42, 43, 44, 45] and boosting both the communication rate and the energy transmission efficiency for simultaneous wireless information and power transfer (SWIPT) as well as wireless powered communications [46, 47, 48, 49, 50, 51, 52].
I-B4 Prototypes/Field Trials of ISs
The versatile promising applications of ISs in wireless networks have spurred extensive worldwide endeavors in experimentation, measurement, and prototyping. Some representative results are summarized in Table I. The usefulness of ISs has been demonstrated in wireless systems operating at frequencies ranging from sub-6 GHz to mmWave, Terahertz (THz), and even optical frequency bands [53, 54, 55, 56].
TABLE I: List of Representative IS Prototypes/Field Trials
Year Description of IS prototypes/field trials
2018
A meta-structure-based reflectarray facilitating 28GHz-band 5G systems was developed by NTT DOCOMO and Metawave [57].
2019
A programmable metasurface-based 8-phase shift-keying transmitter operating at the 4.25 GHz band was presented [58].
2020
A surface with over 3200 elements operating at the 2.4 GHz band was developed, improving signal strength [59].
2020
A 36-element array of inexpensive antennas operating at the 2.4 GHz band was developed, improving channel capacity [60].
2020
A high-gain yet low-cost RIS was developed with 256 elements and 2-bit phase shifting operating at the 2.3 GHz or 28.5 GHz band [61].
2020
A smart surface prototype was developed with measurements showing throughput improvement over commercial MIMO systems [62].
2021
An RIS path loss model was verified by experimental measurements conducted at the 4.25 GHz and the 10.5 GHz band [63].
2021
An RIS prototype with 1100 elements working at the 5.8 GHz band was developed, achieving high performance in various scenarios [64].
2022
An IRS prototype operating at the 2.6 GHz band was developed by Huawei and its collaborators, showing superiority to 5G systems [65].
2022
A prototype of IOS-based wireless communications was developed with the functionality of IOS verified by experiments [66].
2022
An RIS prototype with the ability to continuously control the phase shifts was developed with its properties characterized [67].
2022
An RF-switch-based RIS prototype operating at the 5.3 GHz band was developed with dataset characterization [68].
2022
RIS prototypes covering sub-6 GHz to 28 GHz bands were developed with field tests conducted by ZTE [69].
2023
Field trials on RIS operating at the sub-6 GHz band were conducted, achieving 40 dB indoor received power improvement [70].
2023
An electronically almost-$360^{\circ}$ steerable metamaterial surface that operates above the 24 GHz band was developed [71].
2023
A multi-RIS prototype was developed to facilitate Wi-Fi and commercial 5G networks operating at the 3.4 GHz and the 5.8 GHz band. [72].
2023
A reconfigurable metasurface operating at the 0.34 THz band was developed, achieving $1^{\circ}$ angular precision of beam scanning [73].
2023
An RIS prototype operating at the 5G FR2 frequency band was developed by Greenerwave and verified by Rohde & Schwarz [74].
2023
Field trials on throughput enhancement with RIS were conducted in a 5G mmWave commercial network at the frequency of 27 GHz [75].
2024
Experiments on using RIS to enhance wireless coverage were conducted in a 5G commercial network operating at the 2.6 GHz band [76].
@@ I-C Standardization Activities
As an innovative radio access technology, IS has already been included in the ITU-R IMT-2030 report on future technology trends, as a candidate technology for enhancing the radio interface of future telecommunication standards [77, 78]. During the last two years, IS has been under evaluation for its inclusion in future telecommunication standards, and pre-standarization and standardization bodies are currently working on it. In this subsection, we briefly overview the most relevant standarization activities on IS. Also, we put them in context with other standarization activities focusing on technologies that have strong ties with IS, notably THz communications and ISAC.
@@ I-C1 Third Generation Partnership Program (3GPP)
  • NR smart repeaters from China Mobile Communications Corporation (CMCC) (RWS-210339) – Study candidate side control information such as timing, beam, and bandwidth information for smart repeaters (RAN1, RAN4); and specification of RF requirements considering side control information (RAN4).
@@
  • NR smart repeaters for Rel-18 from Qualcomm (RWS-210019) – Work in Rel-18 on smart repeaters, including the study of side control information (RAN1), protocol support (RAN2), and RF and electromagnetic compatibility requirements (RAN4).
@@
As far as 3GPP is concerned, Rel-18 will include, on the other hand, a study item (RP-213700 [80], TR-38.867 [81]) on network controlled repeaters (NCRs). These repeaters are intended to address the limitations of conventional RF repeaters as well as to provide enhanced low-cost coverage. Conventional amplify-and-forward repeaters are considered which operate with minimal processing and do not require signal decoding and reconstruction. Therefore, they operate at a lower complexity and latency than decode-and-forward repeaters. The price to pay for the reduction in complexity lies in a reduced link quality and increased network interference, which is due to the interference/noise amplification. When applied to high frequency bands (above 24 GHz), amplify-and-forward repeaters provide limited coverage. In fact, static directional repeaters cannot adapt to the channel conditions and user mobility, which is necessary in communication systems where directional communications and high beamforming gains are necessary. NCRs will be designed to include an in-band control interface to configure and manage their behavior, with the objective of reducing the amplification of noise and supporting communications with a better spatial directivity. A detailed description of NCRs in the context of 3GPP standardization can be found in [82]. NCRs are often considered a stepping stone for ISs, since the control channel may be reused. Recent research works have also compared NCRs and ISs, as summarized in [82].
@@ I-C2 European Telecommunication Standards Institute (ETSI)
TABLE II: List of Representative Overview/Survey/Tutorial Papers on ISs
@@
Communication Models
Offer a comprehensive tutorial on electromagnetically consistent communication models for IS, including discrete-type and continuous-type IS.
[90]
2022
Fundamentals of Metamaterial-based IS
Offer a comprehensive tutorial on the theory behind metamaterial-based RIS, including near-field channel models.
[91]
2020
@@
Provide an overview of RIS and explore the theoretical performance limits of RIS-aided wireless communications
@@
2019
Provide a comprehensive overview of signal processing techniques on RIS/IRS-aided wireless systems.
[94]
2022
@@
Provide an overview on using IS for low-complexity data modulation.
[100]
2021
@@
Discuss the concept, applications, challenges and future research directions of RIS/IRS.
[102]
2021
@@ I-D Aims and Organization
In contrast to the existing overview, survey, and tutorial papers on ISs listed in Table I-C2 and two recent books [113, 114], this paper aims to provide a comprehensive overview of the latest development and advancement of ISs, with an emphasis on recent research results as well as new IS architectures and their design issues. Specifically, this paper first provides an in-depth overview of the state-of-the-art results on the design of the commonly adopted reflection-based IS including its reflection optimization, deployment, signal modulation, wireless sensing, and ISAC. Next, we delve into the recent progress in advanced IS architectures and their associated new design issues and solutions. Moreover, we highlight the important problems that remain unsolved to motivate future research on ISs. We hope that this paper will help unveil the immense potential of ISs for applications in future wireless networks such as 6G, and provide useful and up-to-date guidance for the future research and development for ISs.